OpenStudy (anonymous):

a triangle has sides of lengths 4, 3, and 5. is it a right triangle? explain.

OpenStudy (anonymous):

yes, it is .

OpenStudy (anonymous):

right triangle holds this propert , H^2 = B^2 + P^2 where H is the longest side

OpenStudy (anonymous):

Hence, 4^2 + 3^2 = 5^2

OpenStudy (anonymous):

ok, so 16=9>25. the legs has to be lower to be a right triangle, don't it?

OpenStudy (anonymous):

is not a right triangle

OpenStudy (anonymous):

It is. since 16 + 9 = 25 !

OpenStudy (anonymous):

oh crap,yeah